A block of Mass M1 rests on a block of mass M2 which lies on a frictionless table.The coeff of friction between the blocks is u.No friction with ground.What is the maximum horizontal force which can be applied to the blocks for them to accelerorce without slipping on one another if the force is applied to a. Block 1,2.Block 2

Grade:12

1 Answers

Chetan Mandayam Nayakar
13 years ago

case1

a=F/(M1+M2),force on M2=M2a =M2F/(M1+M2)=uM1g, therefore Fmax=uM1(M1+M2)g/M2

case2

a=F/(M1+M2), force on M1= M1a =M1F/(M1+M2)=uM1g, therefore Fmax= u(M1+M2)g

'Win exciting gifts by answering the questions on Discussion Forum.'